# A nonequilibrium quantum Otto engine enhanced via multi-parameter control

### v1

@@ IV Two-parameter enhancement
$-\Delta W$ $=\frac{1}{2}(g_{h}-g_{l})\left(\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}-\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}\right)$ (19)
@@
This expression enables a direct comparison between the net work of the single-parameter Otto cycles with the sub-cycles of the two-parameter Otto cycle, $W^{g}$ and $W^{\omega}$, where $W\!=\!W^{g}\!+\!W^{\omega}$, as defined after Eq. (III).
@@
To proceed with such a comparison, we note that an increase to the harmonic trapping frequency, $\omega_{l}\!\to\!\omega_{h}$ with $\omega_{h}\!>\!\omega_{l}$, for fixed interaction strength $g_{h}$, compresses the atomic cloud, thereby increasing the integral of the squared density profile, $\int dx\,\rho(x)^{2}$. To guarantee an overall increase to the integrated correlation function, given below Eq. (8) as $\langle\hat{\overline{G_{2}}}\rangle\!=\!\int dx\,g^{(2)}(x)\rho(x)^{2}$, we additionally require that $g^{(2)}(x)$ grows under this same compression. Indeed, we know that $g^{(2)}(x)$ monotonically decreases with the dimensionless interaction strength $\gamma(x)\!=\!mg/\hbar^{2}\rho(x)$ [67]. As this dimensionless interaction strength is inversely proportional to the density, we find that the total integrated correlation function indeed increases with the trapping frequency, meaning $\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}\!>\!\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}$ for $\omega_{h}\!>\!\omega_{l}$. Therefore, the net work extracted from the interaction-driven sub-cycle of the two-parameter Otto cycle, given on the first line of Eq. (IV), exceeds that extracted from the interaction-driven single-parameter cycle in Eq. (17).
@@
Similarly, increasing the interaction strength, $g_{l}\!\to\!g_{h}$ with $g_{h}\!>\!g_{l}$, at a fixed value of the harmonic trapping frequency, $\omega_{h}$, broadens the atomic cloud, as increased interparticle repulsion drives the atoms further apart. This results in an increased second moment of the density distribution, $\langle x^{2}\rangle$, meaning $\langle x^{2}\rangle_{(g_{h},\omega_{h})}\!>\!\langle x^{2}\rangle_{(g_{l},\omega_{h})}$. From this, we find that the contribution to $W_{\mathbf{C}\to\mathbf{D}}$ from the volumetric sub-cycle of the two-parameter Otto cycle, given on the second line of Eq. (IV), exceeds $W_{\mathbf{C}\to\mathbf{D}}^{\omega}$ for the single-parameter Otto cycle, given in Eq. (18).
@@
Therefore, by comparing both terms contributing to the two-parameter Otto cycle against their single-parameter counterparts, we arrive at the conclusion that $W_{\mathbf{C}\to\mathbf{D}}\!>\!-W_{\mathbf{C}\to\mathbf{D}}^{g}\!-\!W_{\mathbf{C}\to\mathbf{D}}^{\omega}$, meaning $\Delta W\!>\!0$. Thus, for the case of a sudden quench of both interaction strength and harmonic trapping frequency in a 1D Bose gas, the net work of the two-parameter quantum Otto cycles is greater than the sum of that for the single-parameter Otto cycles taken in isolation. We additionally note that the same arguments apply to any Otto cycle where control is over both interaction strength and volume of the gas. In particular, the arguments presented above also apply to the the 1D Bose gas confined to either a uniform box trap or within a ring, where the total system length is controllable. Such a scenario may be thought of as a quantum analogue of the classical piston engine.
More generally, one may consider whether there is a benefit to operating any two-parameter sudden quench Otto cycle over the related single-parameter cycles. In particular, in Appendix B, we examine an Otto cycle where the sudden quench is taken for the scalar strength parameters $c^{(1)}$ and $c^{(2)}$, as first introduced for a general physical model in Eq. (1), and which are associated with the operators $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$, respectively. The net work of this sudden quench Otto cycle is given by Eq. (4), for a sudden quench over the parameters $c^{(1)}$ and $c^{(2)}$ between their respective values in the low ($l$) and high ($h$) energy equilibrium states. Following arguments similar to those laid out above for the case of a 1D Bose gas, we find that the net work of the two-parameter Otto cycle out-performs the sum of the single-parameter Otto cycles if both $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$ grow under an increase to their complementary strength parameters, i.e. $c^{(2)}$ and $c^{(1)}$, respectively.
An application for such an inequality, beyond the 1D Bose gas already discussed, is the harmonically trapped 1D Fermi gas, commonly known as the Yang-Gaudin model [69, 70]. There, one may additionally utilize control over the spin polarization to enhance performance, as recently explored for a uniform system in Ref. [21]. Further, one may consider applying a two-parameter quench in the context of the transverse-field Ising model, where it is possible to exploit control over both the nearest-neighbor interaction and the transverse magnetic field [71, 72, 73]. While application to these models would be of interest, in the following we restrict our attention to the case of a harmonically trapped 1D Bose gas.
@@
Finally, we note that the enhanced performance of a two-parameter quench thermal machine over its single-parameter constituents, expressed as $\Delta W>0$, remains true regardless of whether the Otto Cycle operates as an engine (which is what we are considering here, in the main text), refrigerator, thermal accelerator, or heater [19], (see Appendix B for further details, where we illustrate this for an Otto refrigerator).
@@ V Zero temperature quasicondensate regime
At zero temperature, the weakly interacting Bose gas may is described by a Thomas-Fermi approximation [62, 61]. From this, one is able to calculate both equilibrium expectation values, $\langle x^{2}\rangle$ and $\langle\hat{\overline{G_{2}}}\rangle$, present in Eq. (III) analytically. More specifically, for the expectation value $\langle x^{2}\rangle$ we obtain $\langle x^{2}\rangle\!=\!(N/5)R_{\mathrm{TF}}^{2}$, where $R_{\mathrm{TF}}\!=\!(3Ng/2m\omega^{2})^{1/3}$ is the Thomas-Fermi radius [61] (for further details, see Appendix C). For the correlation function $\langle\hat{\overline{G_{2}}}\rangle$, on the other hand, we obtain the approximation $\langle\hat{\overline{G_{2}}}\rangle\simeq\!g^{(2)}(0)\int dx\rho(x)^{2}\!\simeq\!bN\rho(0)g^{(2)}(0)$, where $\rho(0)\!=\!(9mN^{2}\omega^{2}/32g)^{1/3}$ is the density of the Thomas-Fermi profile at the trap center, and $b$ is a constant factor determined by the form of the density profile, with b=4/5b\!=\!4/5 for the Thomas-Fermi inverted parabola. Thus, the integrated correlation function depends on
@@
Combining these approximations in Eqs. (5) and (III), we thus obtain simple analytic results for the efficiency and net work of the two-parameter sudden quench chemical Otto engine and illustrate them in Fig. 2, as functions of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, the system is initialized in the low-energy equilibrium state with N=2000N\!=\!2000 particles. Then, following the work input stroke, $W_{\mathbf{A}\to\mathbf{B}}$, the working fluid takes in $\Delta N\!=\!200$ particles during the equilibration stroke $\mathbf{B}\!\to\!\mathbf{C}$ with the high energy reservoir. To operate in a closed cycle, the same number of particles is later output into the low energy reservoir during the corresponding equilibration stroke $\mathbf{D}\!\to\!\mathbf{A}$ [78].
@@
Figure 3: Comparison of the two-parameter sudden quench quantum Otto engine cycle against the single-parameter Otto engine cycles. In panel (a), we demonstrate the net work of the single-parameter Otto engine cycles as a function of their quenched parameter; the interaction-driven cycle, $W_{g}$, is shown as the solid blue line, whereas the volumetric cycle, $W_{\omega}$, as the solid red line. The value of the quenched parameter ratio is denoted ‘ratio’. This is contrasted with the maximum net work of the two-parameter Otto engine cycle, denoted $W^{\mathrm{max}}$ and shown as the solid yellow line. The maximum net work of both single-parameter Otto cycles are given by the blue and red dashed lines for the interaction-driven and volumetric cycles, respectively, with their sum shown as the black dashed line. In contrast, the maximum net work of the two-parameter Otto engine cycle is greater than this sum by more than an order of magnitude. Panel (b) demonstrates the efficiency of the single-parameter Otto engine cycles, with colors corresponding to those shown in panel (a). The efficiency at maximum net work of the two-parameter Otto cycle is shown as the solid yellow line, and clearly out-performs both single-parameter cycles both in terms of magnitude and breadth of operation.
@@
Previously, in Sec. IV, we have shown that, for a working fluid consisting of a harmonically trapped 1D Bose gas, the performance of the two-parameter Otto cycle generally out-performs the sum of the individual cycles taken in isolation. Here, in this concrete example of a two-parameter Otto cycle, we see that this engine greatly outperforms the individual Otto cycles, as the two-parameter cycle generates net work that is more than an order of magnitude greater than the sum of both individual engine cycles operating at their respective maxima (shown as the black dashed line in Fig. 3(a)). Additionally, though it was not guaranteed by the analysis presented in Sec. IV, we find that the efficiency of the two-parameter Otto engine cycle also outperforms what is achieved by both single-parameter engine cycles.
@@
The additional terms contributing to the net work, given in Eq. (25), contain terms arising from the corrections to the coherent correlations, i.e. from $g^{(2)}(0)\!-\!1\!=\!2\sqrt{\gamma(0)}/\pi$. It is these corrections to the correlation function that restrict engine operation to a limited range of the parameter ratios. In particular, the dimensionless interaction strength of the high energy equilibrium state, denoted $\gamma_{0,h}$, scales as $\gamma_{0,h}\!\propto\!r$, causing this contribution to reduce the overall net work. At large enough values of this ratio, these correction terms result in W>0W\!>\!0, meaning the cycle no longer operates as an engine. We therefore note that taking account of these corrections to coherent correlations is essential to correctly evaluating the performance of this quantum Otto cycle.
VI Finite temperature quasicondensate
@@
At low but finite temperatures such that $2\gamma_{0}\!\ll\!\tau_{0}\!\ll\!2\sqrt{\gamma_{0}}$, the weakly interacting ($\gamma_{0}\!\ll\!1$) 1D Bose gas inhabits the thermal quasicondensate regime [67]. Here, the Thomas-Fermi approximation for the density profile is no longer a good approximation, and must be replaced by the thermodynamic Bethe ansatz solution under a local density approximation [53, 66]. This method is numerically exact, and enables calculation of both equilibrium expectation values required to evaluate engine performance via Eqs. (5) and (III). At non-zero temperature, we may additionally utilize a finite temperature difference between the two equilibrium states $\mathbf{A}$ and $\mathbf{C}$ (see Appendix A), meaning our working fluid exchanges both heat and particles with the reservoirs during equilibration, i.e. we investigate operation of a thermo-chemical Otto engine cycle [20].
@@
Figure 4: Performance of the two-parameter sudden quench thermo-chemical quantum Otto engine cycle in the finite temperature quasicondensate regime, evaluated via numerically exact TBA methods. Panels (a) and (b) demonstrate the net work and efficiency, respectively, as a function of the ratio of both quenched parameters. The parameter values for the low energy equilibrium state are chosen to match those used in Fig. 2, but at a finite dimensionless temperature of $\tau_{0}\!=\!1.2\!\times\!10^{-2}$ (see text). Further, we utilize a fixed temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$, making this a thermo-chemical quantum Otto cycle. We observe a maximum net work of $W/\hbar\omega_{l}\!\simeq\!1.1\!\times\!10^{4}$, corresponding to $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$, and an efficiency at maximum net work of $\eta\!\simeq\!0.04$. Panels (c) and (d) display a comparison between the maximum net work and efficiency at maximum net work of the two parameter Otto cycle, respectively, against the single-parameter Otto cycles, as previously shown for the zero temperature quasicondensate system in Fig. 3.
@@
Performance of the two-parameter thermo-chemical quantum Otto cycle, in terms of net work and efficiency, for a harmonically trapped 1D Bose gas in the quasicondensate regime is shown in Fig. 4 (a) and (b). Here, we use the same total net particle exchange of $\Delta N\!=\!200$ as in Fig. 2, but with an additional temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$. We note that, in Ref. [19], it was found that—for a harmonically trapped system (as opposed to the uniform case)—it was essential to enable diffusive contact and chemical work in order that net beneficial work could be extracted from the interaction-driven Otto cycle under a sudden quench.
@@
We observe that, when compared to the pure chemical quantum Otto cycle, operating at $T=0$ and shown in Fig. 2, the two-parameter thermo-chemical quantum Otto cycle investigated here operates as an engine for a broader range of both quenched parameter ratios. This difference stems from the fact that we are incorporating a finite temperature difference between the high and low energy equilibrium states, in addition to the particle number difference. This increase to temperature in the high energy equilibrium state broadens the atomic density profile, increasing the second moment of the density distribution $\langle x^{2}\rangle$, and hence resulting in an enhancement to the net work extracted from the volumetric sub-cycle (see Eq. (III)).
@@
In panels (c) and (d) of Fig. 4, we contrast the maximum work and efficiency at maximum work of the two-parameter Otto cycle against the performance of the single-parameter Otto cycles in the finite-temperature quasicondensate regime. Through this, we observe a significant enhancement to the performance when utilizing a two-parameter quench, similar to what was observed in zero-temperature quasicondensate system in Fig. 3. Notably, the maximum value of net work achieved from this two-parameter Otto engine cycle is again approximately an order of magnitude greater than the sum of that achieved by the single-parameter Otto engine cycles, shown as the black dashed line in Fig.  4  (c), with a similar improvement to the efficiency at maximum work.

### v2

@@ IV.1 Two-parameter enhancement
$-\Delta W$ $=\frac{1}{2}(g_{h}-g_{l})\left(\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}-\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}\right)$
@@
This expression enables a direct comparison between the net work of the single-parameter Otto engines with the sub-cycles of the two-parameter Otto engine, $W^{g}$ and $W^{\omega}$, where $W\!=\!W^{g}\!+\!W^{\omega}$, as defined after Eq. (IV).
@@
To this end, we first note that any increase to the harmonic trapping frequency, $\omega_{l}\!\to\!\omega_{h}$ with $\omega_{h}\!>\!\omega_{l}$, for fixed interaction strength $g_{h}$, compresses the atomic cloud, thereby increasing the integral of the squared density profile, $\int dx\,\rho(x)^{2}$. To guarantee an overall increase to the integrated correlation function, given below Eq. (22) as $\langle\hat{\overline{G_{2}}}\rangle\!=\!\int dx\,g^{(2)}(x)\rho(x)^{2}$, we additionally require that $g^{(2)}(x)$ grows under this same compression. Indeed, we know that $g^{(2)}(x)$ monotonically decreases with the dimensionless interaction strength $\gamma(x)\!=\!mg/\hbar^{2}\rho(x)$ [70]. As this dimensionless interaction strength is inversely proportional to the density, we find that the total integrated correlation function indeed increases with the trapping frequency, meaning $\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}\!>\!\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}$ for $\omega_{h}\!>\!\omega_{l}$.
@@
Similarly, increasing the interaction strength, $g_{l}\!\to\!g_{h}$ with $g_{h}\!>\!g_{l}$, at a fixed value of the harmonic trapping frequency, $\omega_{h}$, broadens the atomic cloud, as increased interparticle repulsion drives the atoms further apart. This results in an increased second moment of the density distribution, $\langle x^{2}\rangle$, meaning $\langle x^{2}\rangle_{(g_{h},\omega_{h})}\!>\!\langle x^{2}\rangle_{(g_{l},\omega_{h})}$.
@@
Thus, for the case of a sudden quench of both interaction strength and harmonic trapping frequency in a 1D Bose gas, the net work of the two-parameter quantum Otto engines is greater than the sum of that for the single-parameter Otto engines taken in isolation. We additionally note that the same arguments apply to any Otto engine where control is over both interaction strength and volume of the gas. In particular, the arguments presented above also apply to the the 1D Bose gas confined to either a uniform box trap or within a ring, where the total system length is controllable. Such a scenario may be thought of as a quantum analogue of the classical piston engine.
@@
We note that the enhanced performance of a two-parameter quench thermal machine over its single-parameter constituents, expressed as $\Delta W\!>\!0$, remains true regardless of whether the Otto cycle operates as an engine (which is what we are considering here, in the main text), refrigerator, thermal accelerator, or heater [19]. We illustrate this for the case of an Otto refrigerator in Appendix B.
@@ V Zero temperature quasicondensate regime
At zero temperature, the weakly interacting Bose gas may is described by a Thomas-Fermi approximation [66, 65]. From this, one is able to calculate both equilibrium expectation values, $\langle x^{2}\rangle$ and $\langle\hat{\overline{G_{2}}}\rangle$, present in Eq. (IV) analytically. More specifically, for the expectation value $\langle x^{2}\rangle$ we obtain $\langle x^{2}\rangle\!=\!(N/5)R_{\mathrm{TF}}^{2}$, where $R_{\mathrm{TF}}\!=\!(3Ng/2m\omega^{2})^{1/3}$ is the Thomas-Fermi radius [65] (for further details, see Appendix C). For the correlation function $\langle\hat{\overline{G_{2}}}\rangle$, on the other hand, we obtain the approximation $\langle\hat{\overline{G_{2}}}\rangle\simeq\!g^{(2)}(0)\int dx\rho(x)^{2}\!\simeq\!bN\rho(0)g^{(2)}(0)$, where $\rho(0)\!=\!(9mN^{2}\omega^{2}/32g)^{1/3}$ is the density of the Thomas-Fermi profile at the trap center, and $b$ is a constant factor determined by the form of the density profile, with b=4/5b\!=\!4/5 for the Thomas-Fermi inverted parabola. Thus, the integrated correlation function depends on
@@
Combining these approximations in Eqs. (5) and (IV), we thus obtain simple analytic results for the efficiency and net work of the two-parameter sudden quench chemical Otto engine and illustrate them in Fig. 1, as functions of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, the system is initialized in the low-energy equilibrium state with N=2000N\!=\!2000 particles. Then, following the work input stroke, $W_{\mathbf{A}\to\mathbf{B}}$, the working fluid takes in $\Delta N\!=\!200$ particles during the equilibration stroke $\mathbf{B}\!\to\!\mathbf{C}$ with the high energy reservoir. To operate in a closed cycle, the same number of particles is later output into the low energy reservoir during the corresponding equilibration stroke $\mathbf{D}\!\to\!\mathbf{A}$ [75].
@@
Figure 2: Comparison of the two-parameter sudden quench quantum Otto engine cycle against the single-parameter Otto engine cycles. In panel (a), we demonstrate the net work of the single-parameter Otto engine cycles as a function of their quenched parameter; the interaction-driven cycle, $W_{g}$, is shown as the solid blue line, whereas the volumetric cycle, $W_{\omega}$, as the solid red line. The value of the quenched parameter ratio is denoted ‘ratio’. This is contrasted with the maximum net work of the two-parameter Otto engine cycle, denoted $W^{\mathrm{max}}$ and shown as the solid yellow line. The maximum net work of both single-parameter Otto engines are given by the blue and red dashed lines for the interaction-driven and volumetric cycles, respectively, with their sum shown as the black dashed line. In contrast, the maximum net work of the two-parameter Otto engine cycle is greater than this sum by more than an order of magnitude. Panel (b) demonstrates the efficiency of the single-parameter Otto engine cycles, with colors corresponding to those shown in panel (a). The efficiency at maximum net work of the two-parameter Otto engine is shown as the solid yellow line, and clearly out-performs both single-parameter cycles both in terms of magnitude and breadth of operation.
@@
Previously, in Sec. IV.1, we have shown that, for a working fluid consisting of a harmonically trapped 1D Bose gas, the performance of the two-parameter Otto cycle generally out-performs the sum of the individual cycles taken in isolation. Here, in this concrete example of a two-parameter Otto engine, we see that this engine greatly outperforms the individual Otto engines, as the two-parameter cycle generates net work that is more than an order of magnitude greater than the sum of both individual engine cycles operating at their respective maxima (shown as the black dashed line in Fig. 2(a)). Additionally, though it was not guaranteed by the analysis presented in Sec. IV.1, we find that the efficiency of the two-parameter Otto engine cycle also outperforms what is achieved by both single-parameter engine cycles.
@@
The additional terms contributing to the net work, given in Eq. (30), contain terms arising from the corrections to the coherent correlations, i.e. from $g^{(2)}(0)\!-\!1\!=\!2\sqrt{\gamma(0)}/\pi$. It is these corrections to the correlation function that restrict engine operation to a limited range of the parameter ratios. In particular, the dimensionless interaction strength of the high energy equilibrium state, denoted $\gamma_{0,h}$, scales as $\gamma_{0,h}\!\propto\!r$, causing this contribution to reduce the overall net work. At large enough values of this ratio, these correction terms result in W>0W\!>\!0, meaning the cycle no longer operates as an engine. We therefore note that taking account of these corrections to coherent correlations is essential to correctly evaluating the performance of this quantum Otto engine.
VI Finite temperature 1D Bose gas
VI.1 Quasicondensate
@@
At low but finite temperatures such that $2\gamma_{0}\!\ll\!\tau_{0}\!\ll\!2\sqrt{\gamma_{0}}$, the weakly interacting ($\gamma_{0}\!\ll\!1$) 1D Bose gas inhabits the thermal quasicondensate regime [70]. Here, the Thomas-Fermi approximation for the density profile is no longer a good approximation, and must be replaced by the thermodynamic Bethe ansatz solution under a local density approximation [57, 54]. This method is numerically exact, and enables calculation of both equilibrium expectation values required to evaluate engine performance via Eqs. (5) and (IV). At non-zero temperature, we may additionally utilize a finite temperature difference between the two equilibrium states $\mathbf{A}$ and $\mathbf{C}$ (see Appendix A), meaning our working fluid exchanges both heat and particles with the reservoirs during equilibration, i.e. we investigate operation of a thermo-chemical Otto engine cycle [20].
@@
Figure 3: Performance of the two-parameter sudden quench thermo-chemical quantum Otto engine cycle in the finite temperature quasicondensate regime, evaluated via numerically exact TBA methods. Panels (a) and (b) demonstrate the net work and efficiency, respectively, as a function of the ratio of both quenched parameters. The parameter values for the low energy equilibrium state are chosen to match those used in Fig. 1, but at a finite dimensionless temperature of $\tau_{0}\!=\!1.2\!\times\!10^{-2}$ (see text). Further, we utilize a fixed temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$, making this a thermo-chemical quantum Otto engine. We observe a maximum net work of $W/\hbar\omega_{l}\!\simeq\!1.1\!\times\!10^{4}$, corresponding to $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$, and an efficiency at maximum net work of $\eta\!\simeq\!0.04$. Panels (c) and (d) display a comparison between the maximum net work and efficiency at maximum net work of the two parameter Otto engine, respectively, against the single-parameter Otto engines, as previously shown for the zero temperature quasicondensate system in Fig. 2.
@@
Performance of the two-parameter thermo-chemical quantum Otto engine, in terms of net work and efficiency, for a harmonically trapped 1D Bose gas in the quasicondensate regime is shown in Fig. 3 (a) and (b). Here, we use the same total net particle exchange of $\Delta N\!=\!200$ as in Fig. 1, but with an additional temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$. We note that, in Ref. [19], it was found that—for a harmonically trapped system (as opposed to the uniform case)—it was essential to enable diffusive contact and chemical work in order that net beneficial work could be extracted from the interaction-driven Otto engine under a sudden quench.
@@
We observe that, when compared to the pure chemical quantum Otto engine, operating at $T=0$ and shown in Fig. 1, the two-parameter thermo-chemical quantum Otto cycle investigated here operates as an engine for a broader range of both quenched parameter ratios. This difference stems from the fact that we are incorporating a finite temperature difference between the high and low energy equilibrium states, in addition to the particle number difference. This increase to temperature in the high energy equilibrium state broadens the atomic density profile, increasing the second moment of the density distribution $\langle x^{2}\rangle$, and hence resulting in an enhancement to the net work extracted from the volumetric sub-cycle (see Eq. (IV)).
@@
In panels (c) and (d) of Fig. 3, we contrast the maximum work and efficiency at maximum work of the two-parameter Otto engine against the performance of the single-parameter Otto engines in the finite-temperature quasicondensate regime. Through this, we observe a significant enhancement to the performance when utilizing a two-parameter quench, similar to what was observed in zero-temperature quasicondensate system in Fig. 2. Notably, the maximum value of net work achieved from this two-parameter Otto engine cycle is again approximately an order of magnitude greater than the sum of that achieved by the single-parameter Otto engine cycles, shown as the black dashed line in Fig.  3  (c), with a similar improvement to the efficiency at maximum work.